OpenStudy (vane11):

I keep getting the wrong answer, so maybe I'm just doing my math wrong, anyway an someone help me out?

OpenStudy (vane11):

A bank wants to get new customers for their credit card. They try two different approaches in their marketing campaign. The first promises a "cash back" reward; the second promises low interest rates. A sample of 500 people is mailed the first brochure; of these, 100 get the credit card. A separate sample of 500 people is mailed the second brochure; 125 get the credit card. Are the two campaigns equally attractive to customers? Compute a 95% confidence interval for the difference in the two proportions. A. (0.21, 0.29) B. (–0.102, 0.002) C. (-0.118, 0.018)

jimthompson5910 (jim_thompson5910):

what did you get for the critical z value?

OpenStudy (vane11):

95% = z* of 1.960

jimthompson5910 (jim_thompson5910):

close enough

jimthompson5910 (jim_thompson5910):

now we need to compute phat1 phat1 = x1/n1 phat1 = 100/500 phat1 = 0.2

jimthompson5910 (jim_thompson5910):

qhat1 = 1 - phat1 qhat1 = 1 - 0.2 qhat1 = 0.8

jimthompson5910 (jim_thompson5910):

the same is done for phat2 and qhat2 phat2 = x2/n2 phat2 = 125/500 phat2 = 0.25 qhat2 = 1 - phat2 qhat2 = 1 - 0.25 qhat2 = 0.75

OpenStudy (vane11):

Ok, I redid it, and I got B, did you get that answer as well?

jimthompson5910 (jim_thompson5910):

Now we use this to calculate the 95% Confidence Interval (CI) Lower Limit: L = (phat1-phat2)-z*sqrt( (phat1*qhat1)/n1 + (phat2*qhat2)/n2 ) L = (0.2-0.25)-1.96*sqrt( (0.2*0.8)/500 + (0.25*0.75)/500 ) L = -0.10167119119973 L = -0.102 ----------------- Upper Limit: U = (phat1-phat2)+z*sqrt( (phat1*qhat1)/n1 + (phat2*qhat2)/n2 ) U = (0.2-0.25)+1.96*sqrt( (0.2*0.8)/500 + (0.25*0.75)/500 ) U = 0.00167119119973 U = 0.002 The 95% confidence interval is (L, U) = ( -0.102, 0.002 )

jimthompson5910 (jim_thompson5910):

yep, got B

OpenStudy (vane11):

Great! Thanks!

jimthompson5910 (jim_thompson5910):

np